# RaGS: Unleashing 3D Gaussian Splatting from 4D Radar and Monocular Cues  for 3D Object Detection

### v2

@@ Abstract
4D millimeter-wave radar has emerged as a promising sensor for autonomous driving, but effective 3D object detection from both 4D radar and monocular images remains a challenge. Existing fusion approaches typically rely on either instance-based proposals or dense BEV grids, which either lack holistic scene understanding or are limited by rigid grid structures. To address these, we propose RaGS, the first framework to leverage 3D Gaussian Splatting (GS) as representation for fusing 4D radar and monocular cues in 3D object detection. 3D GS naturally suits 3D object detection by modeling the scene as a field of Gaussians, dynamically allocating resources on foreground objects and providing a flexible, resource-efficient solution. RaGS uses a cascaded pipeline to construct and refine the Gaussian field. It starts with the Frustum-based Localization Initiation (FLI), which unprojects foreground pixels to initialize coarse 3D Gaussians positions. Then, the Iterative Multimodal Aggregation (IMA) fuses semantics and geometry, refining the limited Gaussians to the regions of interest. Finally, the Multi-level Gaussian Fusion (MGF) renders the Gaussians into multi-level BEV features for 3D object detection. By dynamically focusing on sparse objects within scenes, RaGS enable object concentrating while offering comprehensive scene perception. Extensive experiments on View-of-Delft, TJ4DRadSet, and OmniHD-Scenes benchmarks demonstrate its state-of-the-art performance. Code will be released.
Introduction
Autonomous driving requires accurate 3D perception for safe decision-making (Mao et al. 2023). Recently, 4D millimeter-wave radar has emerged as a highly promising sensor, known for its robustness in challenging environments, long-range detection capabilities, and ability to capture both velocity and elevation data. Cameras, on the other hand, provide high-resolution semantic information. The complementary nature of these two modalities is crucial for enhancing perception in autonomous driving, particularly for 3D object detecting. However, effective as well as efficient fusion across modalities remains a challenge, as the task is inherently sparse while still requiring a comprehensive understanding of the entire scene.
@@ Introduction
Figure 2: Pipeline of RaGS. RaGS consists of a Feature Extractor & Head, Frustum-based Localization Initiation (FLI), Iterative Multimodal Aggregation (IMA), and Multi-level Gaussian Fusion (MGF). The positions of the Gaussians are initialized using the FLI module, along with learnable attributes such as rotation, scale, opacity, and implicit feature embeddings. These Gaussians are then passed into the IMA module, where they are projected onto the image plane to gather semantic information. Next, they are processed as voxels using sparse convolution with height-extended radar geometry. Residuals relative to regions of interest are computed iteratively, updating the positions towards sparse objects. Finally, the multi-level Gaussians are rendered into Bird’s Eye View (BEV) features and fused through MGF, followed by cross-modal fusion for 3D object detection.
@@
Recent research has focused on either instance-based or BEV-based fusion for 4D radar and camera 3D object detection. Instance-based approaches (Kim et al. 2023a; Long et al. 2023; Nabati and Qi 2021) rely on 2D detectors to generate proposals, which are refined by aligning radar features. While these approaches directly maintain the 3D detection flow by focusing on sparse objects, they lack global scene understanding and are constrained by cascaded network designs, as shown in Fig. 1(a). On the other hand, BEV-based approaches (Kim et al. 2023b; Xiong et al. 2024; Zheng et al. 2023; Lin et al. 2024; Bai et al. 2024) project multimodal features into a predefined, fixed top-down space, enabling global reasoning. However, these approaches suffer from rigid voxelization, fixed anchor positions for sampling image semantics, and inefficiencies due to excessive background aggregation, as illustrated in Fig. 1(b). While they offer promising performance in building comprehensive scene perception, they are misaligned with the sparse nature of 3D object detection tasks, in which focusing primarily on foreground objects is significantly more efficient.
@@
In parallel, 3D Gaussian Splatting (GS) (Kerbl et al. 2023) has emerged as a compact, continuous scene representation. Originally developed for neural rendering, GS models scenes as collections of anisotropic Gaussians optimized in 3D space, offering physical interpretability, sparsity, and flexibility. Recent extensions have applied GS to dynamic scenes (Yang et al. 2024) and LiDAR-camera fusion (Hess et al. 2025). However, these works have primarily focused on rendering or occupancy prediction (Huang et al. 2024), without exploring its potential for multi-modal fusion or 3D object detection, despite its natural alignment with the flexible needs of such tasks. Unlike 2D instance-based or BEV-based approaches, GS can dynamically allocate resources and adjust its attention, making it particularly efficient for identify interested objects within entire scene.
@@
To address the limitations of existing fusion approaches and better adapt to the sparse nature of 3D object detection, we propose RaGS, a novel framework that leverages 3D Gaussian Splatting to fuse 4D radar and monocular cues, as presented in Fig. 1(c). Unlike traditional dense voxel grids, RaGS represents the scene as a continuous field of 3D Gaussians, allowing for flexible aggregation without the constraints of fixed grid resolutions and fixed anchor sampling. By dynamically allocating resources iteratively, RaGS focuses on sparse foreground objects while preserving global scene perception through Gaussians’s continuous probability distribution. Specifically, we follow a cascaded pipeline composed of three key modules, as presented in Fig. 2. First, the Frustum-based Localization Initiation (FLI) identifies foreground regions and unprojecting pixels to initialize coarse, grounded 3D Gaussian localization. Next, the Iterative Multimodal Aggregation (IMA) aggregates radar and image features within frustum regions, refining Gaussian positions towards objects and enhancing both semantic and geometric expressiveness. Finally, the Multi-level Gaussian Fusion (MGF) integrates features rendered from Gaussians at different levels to produce a rich representation for object prediction. We evaluate RaGS on three 4D radar and camera benchmarks: View-of-Delft (VoD) (Palffy et al. 2022), TJ4DRadSet(Zheng et al. 2022), and OmniHD-Scenes(Zheng et al. 2024), achieving state-of-the-art performance across the datasets, demonstrating the effectiveness of continuous Gaussian representation for multi-modal 3D object detection. Our contributions are summarized as
@@
Related Work
4D Radar based 3D Object Detection
@@
Recent work on 3D object detection using 4D radar and camera fusion typically follows either instance-based or BEV-based paradigms. Instance-based approaches generate 2D proposals from images, refined with radar features. Examples include CenterFusion (Nabati and Qi 2021) that aligns radar points with 2D boxes for better center predictions, CRAFT (Kim et al. 2023a) that integrates radar range and velocity cues, and RADIANT (Long et al. 2023) that recovers missed detections under occlusion. While effective for foreground localization, these approaches depend heavily on proposal quality and lack global perception.
@@
BEV-based approaches project multi-modal features into a top-down space for global spatial reasoning. RCFusion (Zheng et al. 2023) utilizes orthographic projection for semantics but neglects depth ambiguity, whereas other approaches like LXL (Xiong et al. 2024) and CRN (Kim et al. 2023b) introduce various improvements for image view transformation. L4DR (Huang et al. 2025b) and V2X-R (Huang et al. 2025a), on the other hand, further explored the role of LiDAR and 4D radar fusion for 3D object detection. Despite their strengths, BEV approaches rely on fixed-resolution voxel grids, leading to inefficiencies, excessive background aggregation, and reduced flexibility in 3D object detection. In contrast, we propose a dynamic 3D Gaussian-based representation, modeling multi-modal features as a continuous Gaussian field. This allows for more flexible and fine-grained fusion, making it particularly suitable for sparse 3D object detection.
3D Gaussian Splatting in Autonomous Driving
@@
3D Gaussian Splatting (GS) has emerged as a key technique for real-time 3D scene reconstruction. Initially introduced as a faster alternative to neural radiance fields for rendering, GS has extended to tasks like non-rigid motion (Yang et al. 2024), large-scale environments (Zhou et al. 2024), and multi-modal fusion with LiDAR-camera (Zhao et al. 2024; Hess et al. 2025) or radar-only inputs (Kung et al. 2025).
@@
Despite these advancements, current GS applications are predominantly in dense prediction tasks, such as rendering or occupancy prediction (Huang et al. 2024). However, its potential for 3D object detection remains largely unexplored, even though the task’s sparse nature inherently aligns with Gaussian flexibility. While NeRF-based approaches (Xu et al. 2023; Tian, Du, and Duan 2023) use volumetric representations for 3D detection, GS offers advantages like faster rendering and better integration with detection pipelines. In this work, we leverage GS for sparse object modeling and scene perception. Unlike traditional BEV grid models that rely on fixed anchors for feature sampling, our method utilizes GS to enable dynamic feature aggregation by adaptively refining Gaussian positions towards targets.
Method
Overview
@@
Feature Extractor
[…]
Then, the metric depth $\mathbf{D}$ is computed by summing over predefined depth bins $\mathbf{D}=\sum_{d=1}^{D}P_{d}\cdot d$, where $P_{d}$ represents the probability of each depth bin $d$. A lightweight segmentation network is further used to predict foreground regions from image features $\mathbf{F}^{\text{2D}}$, producing segmentation logits $\mathbf{L}\in\mathbb{R}^{H\times W}$. For the 4D radar point cloud, to fully utilize its RCS and velocity characteristics, we pillarize it and apply RCS-Vel-Aware sparse convolution to extract channel-specific information followed (Zheng et al. 2023). These channel-splited features are then stacked into sparse pillars $\mathbf{F}^{\text{pillar}}\in\mathbb{R}^{G\times C}$, where $G$ is the number of pillars. This separate extraction helps avoid confusion and guides the subsequent feature aggregation of the Gaussian.
Gaussian Aggregation
@@
[…]
@@
where $\mathbf{K}$ is the camera intrinsic matrix. Furthermore, we also include 4D radar points $\mathbf{P}_{\text{radar}}$ and randomly sample points in the frustum space to enrich spatial coverage and eliminate potential instability of foreground identification. For the random sampling process, we begin by defining the candidate points $\mathbf{P}_{\text{cand}}$ as the centers of predefined voxels in the radar coordinate system. These points are projected onto the image using the camera’s extrinsic parameters:
$(ud,vd,d)=\mathbf{K}\left(\mathbf{R}\mathbf{P}_{\text{cand}}+\mathbf{T}\right),$ (2)
where $\mathbf{R}$ and $\mathbf{T}$ are the rotation matrix and translation vector from radar to camera coordinates. We select points that lie within the image boundaries using the condition $\mathbf{I}=\{i\mid 0\leq u_{i}<W,0\leq v_{i}<H\}$, and apply Furthest Sampling:
@@
The final localization $\mathbf{P}$ is a concatenation of the unprojected foreground points, 4D radar points, and sampled points:
$\mathbf{P}=\mathtt{Concat}(\mathbf{P}_{\text{unproj}},\mathbf{P}_{\text{sample}},\mathbf{P}_{\text{radar}})\in\mathbb{R}^{N\times 3}.$ (4)
This ensures that all initialized Gaussians are within the field of view (FoV) and offer coarse-but-structured initial localization, unlike BEV-based approaches that perceive the entire scene. The fully learnable explicit features include a rotation vector $\mathbf{R}\in\mathbb{R}^{N\times 4}$, scale $\mathbf{S}\in\mathbb{R}^{N\times 3}$, and opacity $\mathbf{O}\in\mathbb{R}^{N\times 1}$, which are further concatenated with the position $\mathbf{P}\in\mathbb{R}^{N\times 3}$ to get final explicit physical attributes $\mathbf{F}^{\text{E}}$:
@@
Here, $N$ is the number of Gaussian anchors we predefined. Additionally, we assign each Gaussian a learnable embedding $\mathbf{F}^{\text{I}}\in\mathbb{R}^{N\times C}$, which serves as a query for aggregating image semantics and radar geometry. As a result, each Gaussian can be represented $\mathbf{G}=\{\mathbf{\mathbf{F}^{\text{E}},\mathbf{F}^{\text{I}}}\}$.
@@
Figure 6: Qualitative comparison with state-of-the-art 4D radar based 3D object detection approaches.
Iterative Multimodal Aggregation (IMA). To enrich the Gaussians with multi-modal information, we iteratively aggregate features from both radar and image domains, which output $M$ Gaussian set, as presented in Fig. 3. Instead of directly projecting Gaussians into 2D image space and applying 2D deformable attention for semantic aggregation, we first construct a depth-aware 3D image feature space out product image feature $\mathbf{{F}^{\text{2D}}}$ with depth probability $\mathbf{{D}^{\text{prob.}}}$. We then perform 3D deformable cross attention (3D-DCA) within it, enabling the Gaussians to interact with semantically and geometrically aligned image features in a spatially consistent manner. Here we treat each Gaussian as a point located at its mean and conduct projection procedure. More specifically, given 3D Gaussian query $\mathbf{F}^{\text{I}}$ at location P, we aggregate image semantics by
@@
where $n$ indexes $T$ sampled offsets around P, $(\textbf{P})$ P ) denotes projection function, $\Delta\textbf{q}\in\mathbb{R}^{3}$ denotes learnable offsets, and $\phi(\cdot)$ performs trilinear interpolation to extract features. The attention weights $\mathbf{A}_{n}\in[0,1]$ and projection matrix $\mathbf{W}$ are learned to guide the aggregation. The utilization of 3D deformable cross-attention (3D-DCA) facilitates coherent interaction between Gaussians in 3D space and image features in the perspective view.
Consequently, we further enhance the implicit feature representation through sparse convolution. Specifically, each Gaussian is treated as a point located at its mean, and the resulting point cloud is voxelized into a sparse grid $\mathbf{V}^{\text{gs}}\in\mathbb{R}^{N\times C}$. Then, instead of relying solely on Gaussian-derived voxels, we additionally incorporate radar pillars $\mathbf{F}^{\text{pillar}}$ to enable synergistic multi-modal learning. Since radar pillars lack vertical resolution, we replicate each pillar along the height dimension to get $\mathbf{V}^{\text{radar}}\in\mathbb{R}^{(G\times Z)\times C}$, here $Z$ is the voxel number along the height dimension, which can be expressed as $\mathbf{V}^{\text{radar}}=\mathtt{Repeat}(\mathbf{F}^{\text{pillar}})$. These radar voxels are inserted into the same voxel grid as the Gaussian-generated ones, and the RCS-velocity aware pillars provide clear physical meaning, thereby better assisting the Gaussians. By concatenating both modalities in a unified voxel space and applying sparse convolution, we treat each gaussian as voxels and enable deep interaction between radar geometry and image semantics. This process is formulated as
@@
[…]
@@
This iterative process, depicted in Fig. 4, progressively refines Gaussian localization for sparse objects, in contrast to the fixed grid sampling of previous BEV-based models, which suffer from inflexibility and inefficiency due to the inevitable excessive background aggregation.
[…]
$\mathbf{F}^{\text{gs}}=\mathtt{Conv}(\mathtt{Concat}(\mathbf{F}^{(1)},\ldots,\mathbf{F}^{(L)})).$ (10)
Subsequently, the fused feature is further enhanced with radar sparse pillars $\mathbf{F}^{\text{pillar}}$ through cross-modal fuser to produce final BEV feature that forwarded to the head for detection, which can be expressed as $\mathbf{F}^{\text{BEV}}=\mathtt{CMF}(\mathbf{F}^{\text{gs}},\mathbf{F}^{\text{pillar}})$. Here, $\mathtt{CMF}$ consists of concatenation and convolution layers. During this process, we render the final-layer Gaussians into a depth map, optimize opacity using LiDAR depth, and use the fused BEV features to predict a segmentation map in the bird’s-eye view, supervised by the occupancy map.
Crucially, while MGF rasterizes to BEV, our Gaussians undergo iterative refinement towards object centric. This fundamentally differs from rigid grid by enabling dynamic aggregation that prioritizes sparse foreground objects, thereby avoiding the background inefficiencies inherent to BEV approaches that rely on fixed anchor sampling. The auxiliary losses further regularize this dynamic process (see supplementary material for details).
Approach Input Entire Annotated Area (%) Driving Corridor (%) FPS
Car Ped Cyclist mAP Car Ped Cyclist mAP
ImVoxelNet (WACV 2022) (Rukhovich 2022) C 19.35 5.62 17.53 14.17 49.52 9.68 28.97 29.39 11.1
RadarPillarNet (IEEE TIM 2023) (Zheng et al. 2023) R 39.30 35.10 63.63 46.01 71.65 42.80 83.14 65.86 98.8
SMURF (IEEE TIV 2023) (Liu et al. 2024) R 42.31 39.09 71.50 50.97 71.74 50.54 86.87 69.72 -
SCKD (AAAI 2025) (Xu et al. 2025) R 41.89 43.51 70.83 52.08 77.54 51.06 86.89 71.80 39.3
FUTR3D (CVPR 2023) (Chen et al. 2023) R+C 46.01 35.11 65.98 49.03 76.98 43.10 86.19 69.32 7.3
BEVFusion (ICRA 2023) (Liu et al. 2023) R+C 37.85 40.96 68.95 49.25 70.21 45.86 89.48 68.52 7.1
RCFusion (IEEE TIM 2023) (Zheng et al. 2023) R+C 41.70 38.95 68.31 49.65 71.87 47.50 88.33 69.23 9.0
RCBEVDet (CVPR 2024) (Lin et al. 2024) R+C 40.63 38.86 70.48 49.99 72.48 49.89 87.01 69.80 -
LXL (IEEE TIV 2023) (Xiong et al. 2024) R+C 42.33 49.48 77.12 56.31 72.18 58.30 88.31 72.93 6.1
LXLv2 (IEEE RAL 2025) (Xiong et al. 2025) R+C 47.81 49.30 77.15 58.09 - - - - 6.5
HGSFusion (AAAI 2025) (Gu et al. 2025) R+C 51.67 52.64 72.58 58.96 88.28 62.61 87.49 79.46 -
SGDet3D (IEEE RAL 2025) (Bai et al. 2024) R+C 53.16 49.98 76.11 59.75 81.13 60.91 90.22 77.42 9.2
RaGS (Ours) R+C 58.15 50.81 76.62 61.86 88.15 61.67 88.69 79.50 10.5
Table 1: Comparison with state-of-the-art approaches on the val set of VoD (Palffy et al. 2022). The best values are in bold, and the second-best values are underlined.
Approach Input $\text{AP}_{\text{3D}}$ (%) $\text{AP}_{\text{BEV}}$ (%)
ImVoxelNet C 14.96 17.12
RadarPillarNet R 30.37 39.24
SMURF R 32.99 40.98
FUTR3D R+C 32.42 37.51
BEVFusion R+C 32.71 41.12
LXL R+C 36.32 41.20
LXLv2 R+C 37.32 42.35
RCFusion R+C 33.85 39.76
HGSFusion R+C 37.21 43.23
SGDet3D R+C 41.82 47.16
RaGS (Ours) R+C 41.72 51.04
Table 2: Comparison with state-of-the-art approaches on the test set of TJ4DRadSet (Zheng et al. 2022).

### v3

@@ Abstract
4D millimeter-wave radar is a promising sensing modality for autonomous driving, yet effective 3D object detection from 4D radar and monocular images remains challenging. Existing fusion approaches either rely on instance proposals lacking global context or dense BEV grids constrained by rigid structures, lacking a flexible and adaptive representation for diverse scenes. To address this, we propose RaGS, the first framework that leverages 3D Gaussian Splatting (GS) to fuse 4D radar and monocular cues for 3D object detection. 3D GS models the scene as a continuous field of Gaussians, enabling dynamic resource allocation to foreground objects while maintaining flexibility and efficiency. Moreover, the velocity dimension of 4D radar provides motion cues that help anchor and refine the spatial distribution of Gaussians. Specifically, RaGS adopts a cascaded pipeline to construct and progressively refine the Gaussian field. It begins with Frustum-based Localization Initiation (FLI), which unprojects foreground pixels to initialize coarse Gaussian centers. Then, Iterative Multimodal Aggregation (IMA) explicitly exploits image semantics and implicitly integrates 4D radar velocity geometry to refine the Gaussians within regions of interest. Finally, Multi-level Gaussian Fusion (MGF) renders the Gaussian field into hierarchical BEV features for 3D object detection. By dynamically focusing on sparse and informative regions, RaGS achieves object-centric precision and comprehensive scene perception. Extensive experiments on View-of-Delft, TJ4DRadSet, and OmniHD-Scenes demonstrate its robustness and SOTA performance. Code will be released.
1 Introduction
@@
Autonomous driving requires accurate 3D perception for safe decision-making [Comprehensive_Survey]. Recently, 4D millimeter-wave radar has emerged as a highly promising sensor, known for its robustness in challenging environments, long-range detection capabilities, and ability to capture both velocity and elevation data. Cameras, on the other hand, provide high-resolution semantic information. The complementary nature of these two modalities is crucial for enhancing perception in autonomous driving, particularly for 3D object detecting. However, effective as well as efficient fusion across modalities remains a challenge, as the task is inherently sparse while still requiring a comprehensive understanding of the entire scene.
@@
Figure 2: Pipeline of RaGS. RaGS consists of a Feature Extractor & Head, Frustum-based Localization Initiation (FLI), Iterative Multimodal Aggregation (IMA), and Multi-level Gaussian Fusion (MGF). The positions of the Gaussians are initialized using the FLI module, along with learnable attributes such as rotation, scale, opacity, and implicit feature embeddings. These Gaussians are then passed into the IMA module, where they are projected onto the image plane to gather semantic information. Next, they are processed as voxels using sparse convolution with height-extended radar geometry, which implicitly utilizes radar velocity to guide residuals movement. Residuals relative to regions of interest are computed iteratively, updating the positions towards sparse objects. Finally, the multi-level Gaussians are rendered into Bird’s Eye View (BEV) features and fused through MGF, followed by cross-modal fusion for 3D object detection.
@@
Recent research has focused on either instance-based or BEV-based fusion for 4D radar and camera 3D object detection. Instance-based approaches [CRAFT, RADIANT, centerfusion] rely on 2D detectors to generate proposals, which are refined by aligning radar features. While these approaches directly maintain the 3D detection flow by focusing on sparse objects, they lack global scene understanding and are constrained by cascaded network designs, as shown in Fig. 1(a). On the other hand, BEV-based approaches [CRN, LXL, RCFusion, RCBEVDet, SGDet3D] project multimodal features into a predefined, fixed top-down space, enabling global reasoning. However, these approaches suffer from rigid voxelization, fixed anchor positions for sampling image semantics, and inefficiencies due to excessive background aggregation, as illustrated in Fig. 1(b). While they offer promising performance in building scene perception, they are misaligned with the sparse nature of 3D object detection tasks, in which focusing primarily on foreground objects is significantly more efficient.
@@
In parallel, 3D Gaussian Splatting (GS) [3DGS] has emerged as a compact, continuous scene representation. Originally developed for neural rendering, GS models scenes as collections of anisotropic Gaussians optimized in 3D space, offering physical interpretability, sparsity, and flexibility. Recent extensions have applied GS to dynamic scenes [DeformableGS] and LiDAR-camera fusion [SplatAD]. However, these works have primarily focused on rendering or occupancy prediction [Gaussianformer], without exploring its potential for multi-modal fusion or 3D object detection, despite its natural alignment with the flexible needs of such tasks. Unlike 2D instance-based or BEV-based approaches, GS can dynamically allocate resources and adjust its attention, making it particularly efficient for identify interested objects within entire scene.
@@
To address the limitations of existing fusion approaches and better adapt to the sparse nature of 3D object detection, we propose RaGS, a novel framework that leverages 3D Gaussian Splatting to fuse 4D radar and monocular cues, as presented in Fig. 1(c). Unlike traditional dense voxel grids, RaGS represents the scene as a continuous field of 3D Gaussians, allowing for flexible aggregation without the constraints of fixed grid resolutions and fixed anchor sampling. By dynamically allocating resources iteratively, RaGS focuses on sparse foreground objects while preserving global scene perception through Gaussians’s continuous probability distribution. Specifically, we follow a cascaded pipeline composed of three key modules, as presented in Fig. 2. First, the Frustum-based Localization Initiation (FLI) identifies foreground regions and unprojecting pixels to initialize coarse, grounded 3D Gaussian localization. Next, the Iterative Multimodal Aggregation (IMA) aggregates radar and image features within frustum regions, refining Gaussian positions towards objects and enhancing both semantic and geometric expressiveness. Finally, the Multi-level Gaussian Fusion (MGF) integrates features rendered from Gaussians at different levels to produce a rich representation for object prediction. We evaluate RaGS on three 4D radar and camera benchmarks: View-of-Delft (VoD) [VoD], TJ4DRadSet [TJ4D], and OmniHD-Scenes [OmniHD], achieving state-of-the-art performance across the datasets, demonstrating the effectiveness of continuous Gaussian representation for multi-modal 3D object detection. Our contributions are summarized as
@@
2 Related Work
2.1 4D Radar based 3D Object Detection
@@ 2 Related Work
In recent years, research on 4D millimeter-wave radar has advanced rapidly, leading to the release of several benchmark datasets and an increasing focus on multi-modal fusion with cameras [TJ4D, VoD, Kradar, OmniHD]. Recent studies on 3D object detection using 4D radar–camera fusion typically follow two paradigms: instance-based and BEV-based paradigms. Instance-based approaches generate 2D proposals from images, refined with radar features. Examples include CenterFusion [centerfusion] that aligns radar points with 2D boxes for better center predictions, CRAFT [CRAFT] that integrates radar range and velocity cues, and RADIANT [RADIANT] that recovers missed detections under occlusion. While effective for foreground localization, these methods depend heavily on proposal quality and lack global perception.
@@
BEV-based approaches project multi-modal features into a top-down space for global spatial reasoning. RCFusion [RCFusion] utilizes orthographic projection for semantics but neglects depth ambiguity, whereas other approaches like LXL [LXL] and CRN [CRN] introduce various improvements for image view transformation. L4DR [L4DR] and Interfusion [InterFusion], on the other hand, further explored the role of LiDAR and 4D radar fusion for 3D object detection. Despite their strengths, BEV approaches rely on fixed-resolution voxel grids, leading to inefficiencies, excessive background aggregation, and reduced flexibility in 3D object detection. In contrast, we propose a dynamic 3D Gaussian-based representation, modeling multi-modal features as a continuous Gaussian field. This allows for more flexible and fine-grained fusion, making it particularly suitable for sparse 3D object detection.
2.2 3D Gaussian Splatting in Autonomous Driving
@@
3D Gaussian Splatting (GS) has emerged as a key technique for real-time 3D scene reconstruction. Initially introduced as a faster alternative to neural radiance fields for rendering, GS has extended to tasks like non-rigid motion [DeformableGS], large-scale environments [Drivinggaussian], and multi-modal fusion with LiDAR-camera [TCLC-GS, SplatAD] or radar-only inputs [RadarSplat].
@@
Despite these advancements, current GS applications are predominantly in dense prediction tasks, such as rendering or occupancy prediction [Gaussianformer]. However, its potential for 3D object detection remains largely unexplored, even though the sparse nature of detection task inherently aligns with Gaussian flexibility. While NeRF-based approaches [Nerf-det, Mononerf] use volumetric representations for 3D detection, GS offers advantages like faster rendering and better integration with detection pipelines.
In this work, we leverage GS for sparse object modeling and scene perception. Unlike traditional BEV grid models that rely on fixed anchors for feature sampling, our method utilizes GS to enable dynamic feature aggregation by adaptively refining Gaussian positions towards targets, which perfectly aligns with the 3DGS foreground fitting insight pioneered by indoor models (3DGS-DET [3DGS-DET], Gaussian-Det [Gaussian-Det]), although RaGS is designed for outdoor environments.
3 Method
@@
3.1 Feature Extractor
The feature extractor processes image and 4D radar. For the monocular image, we employ a ResNet-50 backbone with an FPN to extract and fuse multi-scale features $\mathbf{C}\in\mathbb{R}^{H\times W\times C}$, which are further fused with sparse radar depth $\mathbf{S}\in\mathbb{R}^{H\times W}$ to estimate depth probability $\mathbf{{D}^{\text{prob.}}}\in\mathbb{R}^{H\times W\times D}$ and enhanced feature map $\mathbf{F}^{\text{2D}}\in\mathbb{R}^{H\times W\times C}$. Here, $C$ and $D$ denotes the number of channel and depth bins, and $(H,W)$ denotes the resolution. The process can be expressed as $(\mathbf{F}^{\text{2D}},\mathbf{{D}^{\text{prob.}}})=\mathtt{Conv}(\mathtt{Concat(\mathbf{C},\mathbf{S})})$, where $\mathtt{Conv}$ denotes convolution that output $(C+D)$ channel feature map.
Then, the metric depth $\mathbf{D}$ is computed by summing over predefined depth bins $\mathbf{D}=\sum_{d=1}^{D}P_{d}\cdot d$, where $P_{d}$ represents the probability of each depth bin $d$. A lightweight segmentation network is further used to predict foreground regions from image features $\mathbf{F}^{\text{2D}}$, producing segmentation logits $\mathbf{L}\in\mathbb{R}^{H\times W}$. For the 4D radar point cloud, to fully utilize its RCS and velocity characteristics, we pillarize it and apply RCS-Vel-Aware sparse convolution to extract channel-specific information followed [RCFusion]. These channel-splited features are then stacked into sparse pillars $\mathbf{F}^{\text{pillar}}\in\mathbb{R}^{G\times C}$, where $G$ is the number of pillars. This separate extraction helps avoid confusion and guides the subsequent feature aggregation of the Gaussian.
3.2 Gaussian Aggregation
@@
Specifically, we begin by selecting the top-K foreground pixels from the segmentation logits $\mathbf{L}$ and metric depth $\mathbf{D}$. Each selected foreground pixel $(u,v)$ with its corresponding depth $d=\textbf{D}(u,v)$ is then unprojected into 3D space:
@@
where $\mathbf{K}$ is the camera intrinsic matrix. Furthermore, we also include 4D radar points $\mathbf{P}_{\text{radar}}$ and randomly sample points in the frustum space to enrich spatial coverage and eliminate potential instability of foreground identification. For the random sampling process, we begin by defining the candidate points $\mathbf{P}_{\text{cand}}$ as the centers of predefined voxels in the radar coordinate system. These points are projected onto the image space using the camera’s extrinsic parameters: $(ud,vd,d)=\mathbf{K}\left(\mathbf{R}\mathbf{P}_{\text{cand}}+\mathbf{T}\right)$, where $\mathbf{R}$ and $\mathbf{T}$ are the rotation matrix and translation vector from radar to camera coordinates. We select points that lie within the image boundaries using the condition $\mathbf{I}=\{i\mid 0\leq u_{i}<W,0\leq v_{i}<H\}$, and apply Furthest Sampling:
@@
The final position $\mathbf{P}$ is a concatenation of the unprojected foreground points, 4D radar points, and sampled points:
$\mathbf{P}=\mathtt{Concat}(\mathbf{P}_{\text{unproj}},\mathbf{P}_{\text{sample}},\mathbf{P}_{\text{radar}})\in\mathbb{R}^{N\times 3},$ (3)
where $N$ is the number of Gaussian anchors. This ensures that all initialized Gaussians are within the field of view (FoV) and offer coarse-but-structured initial localization, unlike BEV-based approaches that perceive the entire scene. The fully learnable explicit properties include a rotation vector $\mathbf{R}\in\mathbb{R}^{N\times 4}$, scale $\mathbf{S}\in\mathbb{R}^{N\times 3}$, and opacity $\mathbf{O}\in\mathbb{R}^{N\times 1}$, which are further concatenated with the position $\mathbf{P}\in\mathbb{R}^{N\times 3}$ to get final explicit physical attributes $\mathbf{F}^{\text{E}}$:
@@
Additionally, we assign each Gaussian with a learnable embedding $\mathbf{F}^{\text{I}}\in\mathbb{R}^{N\times C}$, which serves as a query for aggregating image semantics and radar geometry. As a result, each Gaussian is represented as $\mathbf{G}=\{\mathbf{\mathbf{F}^{\text{E}},\mathbf{F}^{\text{I}}}\}$.
@@
Iterative Multimodal Aggregation (IMA). To enrich the Gaussians with multi-modal information and progressively guide them toward the foreground regions, we iteratively aggregate features from both image and radar domains. In particular, the spatial cues from radar serve as anchors for residual position updates, while the velocity features provide implicit guidance for gradual convergence. This iterative process outputs the final set of $M$ refined Gaussians, as illustrated in Fig. 3.
For semantic aggregation, instead of directly projecting Gaussians into 2D image space and applying 2D deformable attention for semantic aggregation, we first construct a depth-aware 3D image feature space out product image feature $\mathbf{{F}^{\text{2D}}}$ with depth probability $\mathbf{{D}^{\text{prob.}}}$. We then perform 3D deformable cross attention (3D-DCA) within it, enabling the Gaussians to interact with semantically and geometrically aligned image features in a spatially consistent manner. Specifically, given 3D Gaussian query $\mathbf{F}^{\text{I}}$ at location P, we aggregate image semantics by
@@
where $n$ indexes $T$ sampled offsets around P, $\mathcal{P}(\textbf{P})$ denotes projection function, $\Delta\textbf{q}\in\mathbb{R}^{3}$ denotes learnable offsets, and $\phi(\cdot)$ performs trilinear interpolation to extract features. The attention weights $\mathbf{A}_{n}\in[0,1]$ and projection matrix $\mathbf{W}$ are learned to guide the aggregation. The utilization of 3D deformable cross-attention (3D-DCA) [SGDet3D] facilitates coherent interaction between Gaussians in 3D space and image features in the perspective view.
Consequently, we further enhance the implicit feature representation using sparse convolution, which integrates radar velocity as motion cues to refine the spatial distribution of Gaussians. Specifically, each optimized Gaussian point is treated as a point located at its center, and the resulting point cloud is voxelized into a sparse grid $\mathbf{V}^{\text{gs}}\in\mathbb{R}^{N\times C}$. Then, we incorporate radar pillars $\mathbf{F}^{\text{pillar}}$ consisting of rich geometry to enable synergistic multi-modal learning. Since radar pillars lack vertical resolution, we replicate each pillar along the height dimension to get $\mathbf{V}^{\text{radar}}\in\mathbb{R}^{(G\times Z)\times C}$, here $Z$ is the voxel number along the height dimension, which can be expressed as $\mathbf{V}^{\text{radar}}=\mathtt{Repeat}(\mathbf{F}^{\text{pillar}})$. These radar voxels are concatenated with $\mathbf{V}^{\text{gs}}$, and the RCS-velocity aware pillars provide clear physical meaning, thereby better assisting the Gaussians. By concatenating both modalities in a unified voxel space and applying sparse convolution, we treat each Gaussian as voxels and enable deep interaction between radar geometry and image semantics. This process is formulated as
@@
where $\mathtt{Spconv}$ denotes sparse convolution. Finally, to maintain consistency with the original Gaussian set, we only output former $N$ voxels using $[:N]$ and reassign to $\mathbf{F}^{\text{I}}$, since the order remains unchanged. After updating the implicit feature, we begin explicit feature refinement as below.
@@
This iterative process, depicted in Fig. 4, progressively refines Gaussian localization for sparse objects, in contrast to the fixed grid sampling of previous BEV-based models, which suffer from inflexibility and inefficiency due to the inevitable excessive background aggregation. All other Gaussian attributes are also updated simultaneously.